# FedMeld: A Model-dispersal Federated Learning Framework for Space-ground Integrated Networks

### v1

@@ I Introduction
Edge intelligence and space-ground integrated networks (SGINs) are two new features of sixth-generation (6G) mobile networks. Edge intelligence aims to deliver pervasive, real-time artificial intelligence (AI) services to users [1, 2], while SGINs ensure ubiquitous connectivity worldwide [3]. The convergence of these two trends in 6G marks a transformative step toward integrating networking and intelligence in space, enabling AI services to reach every corner of the planet. For instance, the Starlink constellation has deployed over 30,000 Linux nodes and more than 6,000 microcontrollers in space [4], creating a satellite server network orbiting around the Earth. This advancement extends traditional 2D edge AI into a 3D space-ground integrated AI paradigm [5], offering mission-critical AI services — such as healthcare, crop monitoring, and emergency response — anytime, anywhere [6]. One primary objective of space-ground integrated AI is to enable federated learning (FL) [5], referred to as space-ground integrated FL (SGI-FL). Based on such technologies, pervasive ground clients collaboratively train AI models without sharing raw data [7, 8], while satellites act as parameter servers to periodically receive from, aggregate over, and redistribute model parameters to ground clients.
@@
SGI-FL generally aims to train a global consensus model across a large-scale geographical area by learning from distributed data, i.e., training a healthcare AI model from hospitals and clinics across sparsely populated areas. Since a parameter server in SGI-FL, such as a low Earth orbit (LEO) satellite, can only cover a limited area [9], cooperation among multiple satellites is often needed. There are two primary approaches to synchronizing models among these satellite servers. The first approach is mixing via ground stations, where each satellite server transmits its aggregated model to a ground station for reaching global consensus. However, since a training round cannot finish until all satellites transmit their local updates to the ground station and receive the updated global model [10, 11], this approach often results in prolonged idle periods of up to several hours per round. To mitigate excessive waiting time, asynchronous FL can be employed, where aggregation is triggered as long as a preset number of models are received by the ground station [12, 13, 14]. Nevertheless, the mixing ratio between current and historical models during weighted averaging in these schemes has been demonstrated only empirically, leaving theoretical analysis unexplored. The other approach is mixing via inter-satellite links (ISLs), where satellites exchange the aggregated parameters with their neighboring satellites through ISLs. This approach can overcome the limited contact time between satellites and ground stations, accelerating the learning process through frequent model exchanges [15, 16]. However, it generates substantial data traffic within the satellite constellations. Furthermore, ISLs are not fully supported in many LEO mega-constellations due to the prohibitively high cost of laser terminals [17]. As a result, many LEO mega-constellations today still rely on bent-pipe frameworks, where data is routed via ground stations [18, 19]. In a nutshell, existing SGI-FL schemes, when serving large-scale regions through multiple satellites, face three major limitations: 1) dependence on ground stations, thereby leading to excessive waiting time, 2) reliance on ISLs, which may not be available, and 3) causing extra communication costs.
SGI-FL features high mobility of satellite servers. Outside the area of SGI-FL, some works have explored mobility-aware FL schemes in terrestrial networks [20, 21, 22, 23]. Such works often adopt discrete Markov chains to model user mobility, using transition probabilities representing the likelihood of mobile users staying with or moving to neighboring edge servers. However, unlike terrestrial users with unpredictable movements, satellites move at high speeds along predetermined and repetitive orbital trajectories [24]. As a result, theoretical models like random walks or Markov chains, commonly used for terrestrial FL, are invalid for SGI-FL. Moreover, in terrestrial networks, parameter servers usually remain stationary. In contrast, in SGI-FL, parameter servers are moving, whereas ground clients are relatively static, as their speeds are negligible as compared with those of satellites [25]. Consequently, existing schemes for mobility-aware terrestrial FL networks cannot be applied to SGI-FL. Effectively supporting FL across large-scale geographical regions in the context of SGINs remains an open problem.
@@
In this paper, to overcome all the above issues, we propose an SGI-FL framework called Federated learning with Model Dispersal (FedMeld) as our answer. FedMeld enables infrastructure-free model aggregation across large-scale geographical regions, eliminating the need for ground stations or ISLs. Our mechanism is inspired by a key observation: Given the repetitive trajectories and store-carry-forward (SCF) capabilities of co-orbital satellites, a satellite can transfer model parameters from one serving region $i$ to the next region $i+1$ by mixing models from these two adjacent areas; Similarly, the circular nature of satellite orbits ensures that the model parameters from region $i+1$ will eventually be returned to region $i$. Consequently, without requiring dedicated ground stations or ISLs, parameter mixing across all regions is naturally achieved through model dispersal via satellite trajectories, akin to how animals disperse seeds across continents. In this way, FedMeld exploits the unique movement patterns of satellites, turning their high mobility from a foe to a friend. Within the FedMeld framework, users can continue training while satellites carrying model parameters traverse between adjacent regions, introducing a latency-accuracy tradeoff in managing model staleness. Moreover, model mixing ratios across regions should also be determined judiciously to achieve optimal learning performance. In this direction, we will answer two fundamental questions in FedMeld: 1) Does the proposed model dispersal method ensure global model convergence? and 2) how can we manage model staleness and mixing ratios to optimize training convergence? Our main contributions are summarized as follows:
@@
  • FedMeld and Convergence Analysis: We propose the FedMeld framework for efficient decentralized FL in SGINs. To support effective control, we theoretically characterize the convergence bound of FedMeld under both full and partial participation scenarios, accounting for the effects of model staleness and model mixing ratio. The results reveal a tradeoff between latency and accuracy. Specifically, increasing the global round interval of model mixing between adjacent regions accelerates the completion of the predefined training time but at the cost of reduced model accuracy. Furthermore, the model mixing ratio exhibits a strong coupling with the degree of model staleness, underscoring the necessity for optimal control strategies.
@@
  • Control of Model Staleness and Mixing Ratio: Building on the theoretical insights, we formulate a joint optimization problem to determine the staleness control and mixing ratio (SC-MR) that minimizes the training loss. This optimization problem is decomposed into two sequential subproblems without compromising optimality. We derive the optimal solutions for both decision variables in closed and semi-closed forms, respectively. The following design insights can be obtained from the theoretical solutions: Under more stringent training latency constraints, the round interval of model mixing becomes larger. Furthermore, when the non-IID degree of data is higher, the mixing ratio of historical models should be reduced.
@@
The remainder of this paper is organized as follows. Models and metrics are introduced in Section II. The FedMeld algorithm and its convergence analysis are detailed in Section III. Building upon the convergence analysis, the joint optimization problem of SC-MR is formulated, and the optimal solutions are derived in Section IV. Experimental results are provided in Section V, followed by concluding remarks in Section VI.
@@ II Models and Metrics
As illustrated in Fig. 1, we consider an FL framework in SGINs, where a mega-LEO satellite constellation collaborates with groups of terrestrial edge devices to train a neural network model. The learning and communication models are described in separate subsections.
Figure 1: An FL system in SGINs.
@@ II-A Learning Model
The constellation consists of multiple circular orbits, with satellites uniformly distributed at equal distance intervals within each orbit. These AI-empowered satellites act as parameter servers and move along predetermined orbits. We assume that there are a set $\mathcal{M}$ areas and a set of $\mathcal{N}$ ground edge devices, with =$\mathcal{M}=\left\{1,\ldots,M\right\}$ and $\mathcal{N}=\left\{1,\ldots,N\right\}$. These ground areas are distributed along a specific orbit, and each area $i\in\mathcal{M}$ covers a subset of devices, denoted by $\mathcal{N}_{i}$ with cardinality $N_{i}$.
@@
The training process consists of $R$ iterations (steps), denoted by =${\mathcal{R}}=\left\{1,\ldots,R\right\}$. We assume that each serving satellite aggregates the model from its covered clients every $E$ iterations, and we define ${\mathcal{I}}_{E}=\left\{{\left.{kE}\right|k=1,2,\ldots,\left\lfloor{\frac{R}{E}}\right\rfloor}\right\}$ as the set of synchronization steps of each area, where $k$ is the index of the global round. Since the devices in the same region are located nearby, we assume that they will be handed over to a new satellite parameter server at the same time according to some criteria such as distance and signal-to-noise ratio (SNR). In addition, considering partial device participation, in the $k$-th global round, the serving (nearest) satellite of area $i$ randomly selects a device set, $\mathcal{U}_{k,i}$ (with cardinality $U_{i}$), out of $\mathcal{N}_{i}$ to participate in the training process.
Let ${\mathbf{w}}_{t,j}$ be the model parameters of device $j$ at step $t\in{\mathcal{R}}$. When $t\in{\mathcal{I}}_{E}$, the models of devices in the same area will be aggregated by the serving satellite. Then, the update rule of device $j$ can be given by
${\mathbf{v}_{t+1,j}}={\mathbf{w}_{t,j}}-{\eta_{t}}{\nabla{F_{j}}\left({\mathbf{w}_{t,j},\varsigma_{t,j}}\right)},$ (1)
where $\eta_{t}$ is the learning rate and $\varsigma_{t,j}$ is the selected batch from device $j$ at step $t$ with $|\varsigma|$ data samples. Here, ${\mathbf{v}_{t+1,j}}$ is an auxiliary variable to denote the intermediate result of one step stochastic gradient descent (SGD) update from ${\mathbf{w}_{t,j}}$.
Motivated by [26], for the full participation scheme, we define two global virtual sequences as ${{\overline{\mathbf{w}}}_{t}}=\frac{1}{M}\sum\limits_{i\in{\mathcal{M}}}{{{\overline{\mathbf{w}}}_{t,i}}}=\frac{1}{M}\sum\limits_{i\in{\mathcal{M}}}{\frac{1}{{{N_{i}}}}\sum\limits_{j\in{{\mathcal{N}}_{i}}}{{{\mathbf{w}}_{t,j}}}}$ and ${{\overline{\mathbf{v}}}_{t}}=\frac{1}{M}\sum\limits_{i\in{\mathcal{M}}}{{{\overline{\mathbf{v}}}_{t,i}}}=\frac{1}{M}\sum\limits_{i\in{\mathcal{M}}}{\frac{1}{{{N_{i}}}}\sum\limits_{j\in{{\mathcal{N}}_{i}}}{{{\mathbf{v}}_{t,j}}}}$, where ${{{\overline{\mathbf{w}}}_{t,i}}}$ and ${{{\overline{\mathbf{v}}}_{t,i}}}$ are area $i$’s virtual sequences. When $t\notin\mathcal{I}_{E}$, both global and area’s sequences are inaccessible. When $t\in\mathcal{I}_{E}$, only ${{{\overline{\mathbf{w}}}_{t,i}}}$ can be fetched for area $i$. After the last training step $R$, we assume that the models can be averaged over all areas to obtain the final global model ${{\overline{\mathbf{w}}}_{R}}$. For the general partial participation scheme, the global virtual sequence $\overline{\mathbf{w}}_{t}$ is substituted by ${\overline{\mathbf{z}}_{t}}=\frac{1}{M}\sum\limits_{i\in{\mathcal{M}}}{\frac{1}{{{U_{i}}}}\sum\limits_{j\in{{\mathcal{U}}_{t,i}}}{{{\mathbf{w}}_{t,j}}}}$.
II-B Communication Model
The training process is conducted in consecutive global rounds until the stopping condition is reached. Since the total latency of each global round is limited, it is reasonable to assume that the network topology of SGINs remains unchanged during each global round. Without loss of generality, we focus on the global round $k$ and one ground area $i\in\mathcal{M}$ for analysis and provide detailed latency expressions as follows.
First, the selected clients conduct $E$ local iterations, and the local computing latency of client $j\in\mathcal{U}_{k,i}$ is calculated as
$T_{j}^{\rm{local}}=\frac{E|\varsigma|\Phi}{L_{j}},\;\forall j\in\mathcal{N},$ (2)
Then, clients in each area upload their models to their connected satellites. We consider orthogonal frequency-division multiple access (OFDMA) for uplink transmissions, and each selected device per global round is assigned a dedicated sub-channel with bandwidth $W^{\rm{U}}$. Then, the uplink rate between device $j$ and its connected satellite at global round $k$ can be expressed as
${C_{k,j}^{\rm{U}}}=W^{\rm{U}}{\log_{2}}\left({1+\frac{{{P_{\rm{UE}}}{G_{\rm{UE}}}{G_{\mathrm{SAT}}}L_{k,j}^{\mathrm{PL}}L^{\mathrm{AL}}}}{{{n_{0}}W^{\rm{U}}}}}\right),$ (3)
where $P_{\rm{UE}}$ is the transmit power of ground devices, $G_{\rm{UE}}$ and $G_{\mathrm{SAT}}$ are transmitting and receiving antenna gain of device and satellite, respectively. The free-space path loss $L_{k,j}^{\mathrm{PL}}$ is defined as $L_{k,j}^{\mathrm{PL}}={\left({\frac{\lambda_{k,j}}{{4\pi{d_{k,j}}}}}\right)^{2}}$, where $\lambda_{k,j}$ is the wavelength of uplink signal and $d_{k,j}$ is the transmission distance. The additional loss $L^{\mathrm{AL}}$ is used to characterize the attenuation due to environmental factors, and $n_{0}$ is the noise power spectral density. After the serving satellite conducts FedAvg with latency $T_{i}^{\rm{agg}}$, the satellite broadcasts the averaged results to the clients, with the downlink rate between satellite and client $j$ being
${C_{k,j}^{\rm{D}}}=W^{\rm{D}}{\log_{2}}\left({1+\frac{{{P_{\rm{SAT}}}{G_{\rm{UE}}}{G_{\mathrm{SAT}}}L_{k,j}^{\mathrm{PL}}L^{\mathrm{AL}}}}{{{n_{0}}W^{\rm{D}}}}}\right),$ (4)
where $W^{\rm{D}}$ is the bandwidth that the satellite broadcasts the global model to the users. In this paper, we focus solely on the applications for users in remote areas. Unlike urban environments with numerous blocks, remote regions have relatively few obstacles. Therefore, we disregard the impact of small-scale fading caused by the multi-path effect [27].
Given the data rate, the transmission latency over uplink and downlink can be expressed as $T_{k,j}^{\mathrm{U}}=\frac{q}{{C_{k,j}^{\text{U}}}}$ and $T_{k,j}^{\mathrm{D}}=\frac{q}{{C_{k,j}^{\text{D}}}}$, where $q$ denotes the number of bits within a model. Then, the total latency of global round $k$ can be given by
$T_{k,i}^{\mathrm{total}}=\mathop{\max}\limits_{j\in{\mathcal{U}_{k,i}}}\left\{{ET_{j}^{\mathrm{local}}+T_{k,j}^{\mathrm{U}}}+T_{k,j}^{\mathrm{D}}\right\}+T_{i}^{\rm{agg}},$ (5)
Since we consider the nearest association strategy, the serving duration of each satellite remains unchanged [28]. Thus, it can be regarded that during the period when a satellite provides connectivity to a specific region, the number of global training rounds it can participate in is fixed, denoted by $K$.
II-C Performance Metric
@@
Here, ${F_{j}}\left(\cdot\right)$ is the local objective, defined by

### v2

@@ I Introduction
Edge intelligence and space-ground integrated networks (SGINs) are two new features of sixth-generation (6G) mobile networks. Edge intelligence aims to deliver pervasive, real-time artificial intelligence (AI) services to users [1, 2]. Satellites are particularly valuable in extending coverage to remote and underserved regions without terrestrial infrastructure, and in serving as a complementary tier to offload traffic in congested urban areas. With their wide coverage footprints enabling successive service across different regions, SGINs can ensure ubiquitous worldwide coverage [3]. The convergence of these two trends in 6G marks a transformative step toward integrating networking and intelligence in space, enabling AI services to reach every corner of the planet. For instance, the Starlink constellation has deployed over 30,000 Linux nodes and more than 6,000 microcontrollers in space [4], creating a satellite server network orbiting around the Earth. This advancement extends traditional 2D edge AI into a 3D space-ground integrated AI paradigm [5], offering mission-critical AI services anytime, anywhere [6].
One primary objective of space-ground integrated AI is to enable federated learning (FL) [5], referred to as space-ground integrated FL (SGI-FL). Based on such technologies, pervasive ground clients collaboratively train AI models without sharing raw data [7, 8, 9], while satellites act as parameter servers to periodically receive from, aggregate over, and redistribute model parameters to ground clients. Recent advances in direct satellite-to-device communication have further strengthened the practicality of this paradigm. With the inclusion of non-terrestrial networks in 3GPP [10, 11] and the commercial adoption of satellite connectivity in mainstream smartphones (e.g., Apple iPhone and Huawei Mate 60), direct user–satellite links have become both technically feasible and cost-effective [12]. A representative use case of SGI-FL is healthcare: hospitals and clinics in sparsely populated areas can jointly train diagnostic or triage models without transmitting sensitive medical records, thereby protecting patient privacy and meeting strict data sovereignty requirements [13, 14]. Similarly, in agriculture and climate monitoring, satellites and dispersed ground sensors can train global predictive models for crop diseases, drought risks, or extreme weather events, where cross-border regulations and site-specific sensitivities make centralized data collection impractical. In both scenarios, SGI-FL enables global-scale model convergence while safeguarding data privacy.
@@
SGI-FL generally aims to train a global consensus model across a large-scale geographical area by learning from distributed data. Since a parameter server in SGI-FL, such as a low Earth orbit (LEO) satellite, can only cover a limited area [15], cooperation among multiple satellites is often needed. Therefore, several classical FL algorithms, such as horizontal FL frameworks [16], are unsuitable for the global training scenarios considered in SGI-FL. These frameworks rely on centralized cloud aggregation, where globally distributed clients are expected to upload model updates directly to a cloud server. Such a design requires stable and persistent connectivity, which is often impractical in SGI-FL schemes [3, 6]. There are two primary approaches to synchronizing models among these satellite servers. The first approach is mixing via ground stations, where each satellite server transmits its aggregated model to a ground station for reaching global consensus. However, since a training round cannot finish until all satellites transmit their local updates to the ground station and receive the updated global model [17, 18], this approach often results in prolonged idle periods of up to several hours per round. The other approach is mixing via inter-satellite links (ISLs), where satellites exchange the aggregated parameters with their neighboring satellites through ISLs. This approach can overcome the limited contact time between satellites and ground stations, accelerating the learning process through frequent model exchanges [19, 20]. However, it generates substantial data traffic within the satellite constellations. Furthermore, ISLs are not fully supported in many LEO mega-constellations due to the prohibitively high cost of laser terminals [21]. As a result, many LEO mega-constellations today still rely on bent-pipe frameworks, where data is routed via ground stations [22, 23]. In a nutshell, existing SGI-FL schemes, when serving large-scale regions through multiple satellites, face three major limitations: 1) global model consensus often results in slow convergence; 2) significant communication costs occur over ISLs in large-scale constellations; and 3) satellite cooperation must rely on extra infrastructure such as ISLs, which are unavailable in many commercial satellites.
@@
In this paper, to overcome all the above issues, we propose an SGI-FL framework called Federated learning with Model Dispersal (FedMeld) as our answer. Our mechanism is inspired by a key observation: Given the repetitive trajectories and store-carry-forward (SCF) capabilities of co-orbital satellites, as shown in Fig. 1, a satellite can transfer model parameters from one serving region $i$ (e.g., A) to the next region $i+1$ (e.g., B) by mixing models from these two adjacent areas; Similarly, the circular nature of satellite orbits ensures that the model parameters from region $i+1$ (e.g., B) will eventually be returned to region $i$ (e.g., A). Consequently, without requiring dedicated ground stations or ISLs, parameter mixing across all regions is naturally achieved through model dispersal via satellite trajectories, akin to how animals disperse seeds across continents. Within the FedMeld framework, users can continue training while satellites carrying model parameters traverse between adjacent regions, introducing a latency-accuracy tradeoff in managing model staleness. Moreover, model mixing ratios across regions should also be determined judiciously to achieve optimal learning performance. The core advantages of FedMeld are summarized as follows:
  • Infrastructure-free model aggregation: By utilizing repetitive trajectories and SCF behaviour of co-orbital satellites, FedMeld eliminates the reliance on ground stations or ISLs for model dispersal across regions, turning the satellite’s orbital mobility from a foe to a friend.
  • Flexible control of latency-accuracy tradeoff: The design introduces tunable parameters for staleness and mixing ratio, enabling clients to keep training without waiting for satellite movement and clients training in other regions.
Figure 1: Illustration of parameter mixing across different regions in the proposed FedMeld framework.
To fully explore the advantages of FedMeld, we will answer two fundamental questions in this paper: 1) Does the proposed model dispersal method ensure global model convergence? and 2) how can we manage model staleness and mixing ratios to optimize training convergence? The key challenges in the convergence analysis of FedMeld primarily lie in the following two aspects. First, while mobility-aware FL schemes have been explored for terrestrial networks [24, 25, 26, 27], these schemes often adopt discrete Markov chains to model user mobility. Unlike terrestrial users, satellites move at high speeds along predetermined and repetitive orbital trajectories [28], resulting in model merging across regions via model dispersal. Therefore, the proposed FedMeld framework differs from existing distributed FL schemes due to its special circular topology. Second, due to computing capabilities and geographical locations, asynchronous FL must be carried out to avoid prolonged client idleness [29, 30, 31, 32], which further complicates convergence analysis. Existing asynchronous FL frameworks, such as FedSat [31], FedAsync [32], and AsyncFLEO [33], address model staleness by introducing heuristic weighting or discounting factors to mix current and outdated models, while FedCM accumulates global gradients across rounds [34]. Although these approaches improve performance empirically, they lack a rigorous theoretical analysis of how staleness impacts convergence. In contrast, our proposed FedMeld provides a theoretical foundation by deriving a semi-closed-form expression for the optimal mixing ratio, which is intuitively justified and validated through simulations. To address these challenges, our main contributions are summarized as follows:
@@
  • FedMeld and Convergence Analysis: We propose the FedMeld framework for efficient decentralized FL in SGINs. To support effective control, we theoretically characterize the convergence bound of FedMeld under both full and partial participation scenarios, explicitly incorporating the impact of satellite constellation, inter-region model staleness, and model mixing ratio. Our analysis reveals a tradeoff between training latency and model accuracy, highlighting the need for carefully designed control strategies.
@@
  • Control of Model Staleness and Mixing Ratio: Building on the theoretical insights, we formulate a joint optimization problem to determine the staleness control and mixing ratio (SC-MR) that minimizes the training loss. This optimization problem is decomposed into two sequential subproblems without compromising optimality. We derive the optimal solutions for both decision variables in closed and semi-closed forms, respectively. The results suggest that tighter training latency leads to higher model staleness, while a higher non-IID degree calls for a lower mixing ratio of historical models.
@@
The remainder of this paper is organized as follows. Models and metrics are introduced in Section II. The FedMeld algorithm and its convergence analysis are detailed in Section III. Building upon the convergence analysis, the joint optimization problem of SC-MR is formulated, and the optimal solutions are derived in Section IV. Experimental results are provided in Section V, followed by concluding remarks in Section VI. The main symbols and parameters used in this paper are summarized in Table I for clarity.
TABLE I: Summary of main notations
Notation Definition
$\mathcal{M},\mathcal{N}$ Set of areas and ground edge devices.
$\mathcal{N}_{i}$, $\mathcal{U}_{k,i}$ Subset of devices in area $i$ and the set of participating clients of area $i$ in the $k$-th global round.
$R$, $E$ Total number of training iterations (steps) and the number of local training iterations performed by each client within the same region before local aggregation by its serving satellite.
$K$ Number of global training rounds that a serving satellite provides for each region.
$\varsigma_{t,j}$, $\eta_{t}$ Selected data batch from device $j$ at step $t$ and the learning rate at step $t$.
${\mathbf{w}}_{t,j}$, ${\mathbf{w}}_{t}$, ${\mathbf{z}}_{t}$ Model parameters of device $j$, global virtual sequence under full participation, and global virtual sequence under partial participation at step $t$.
$F\left(\mathbf{w}\right)$, $F_{j}\left(\mathbf{w}\right)$ Global loss function and local loss function of client $j$.
$\delta$, $\alpha$ Global round interval and mixing ratio of model mixing between adjacent regions.
$T_{k,i}^{{\rm{total}}}$, $T_{i,i+1}^{{\rm{idle}}}$, $T_{i,i+1}^{{\rm{fly}}}$ Total training latency of area $i$ in global round $k$, total duration from the SCF satellite leaving area $i$ until completing model aggregation for area $\left(i+1\right)$, and idle duration of clients in the region $(i+1)$.
$T_{\max}$ Maximum tolerable training time.
$L,\mu,G,\Gamma$ Hyperparameter related to assumptions in the convergence analysis.
$f\left(\delta,\alpha\right)$ Convergence bound of FedMeld as a function of $\delta$ and $\alpha$.
$\zeta_{1},\zeta_{2},\zeta_{3},\kappa_{1},\kappa_{2}$ Key parameters related to the convergence bound of FedMeld.
@@ II Models and Metrics
We consider an FL framework in SGINs, where a mega-LEO satellite constellation collaborates with groups of terrestrial edge devices to train a neural network. The constellation consists of multiple circular orbits, with satellites uniformly distributed at equal distance intervals within each orbit. These AI-empowered satellites act as parameter servers and move along predetermined orbits. We assume that there are a set $\mathcal{M}$ of areas and a set $\mathcal{N}$ of ground edge devices, with $\mathcal{M}=\left\{1,\ldots,M\right\}$ and $\mathcal{N}=\left\{1,\ldots,N\right\}$. These ground areas are distributed along a specific orbit, and each area $i\in\mathcal{M}$ covers a subset of devices, denoted by $\mathcal{N}_{i}$ with cardinality $N_{i}$. The learning and communication models are described as follows.
@@ II-A Learning Model
We first introduce the procedure of FedAvg underpinning FedMeld. The training process consists of $R$ iterations (steps), denoted by ${\mathcal{R}}=\left\{1,\ldots,R\right\}$. We assume that each serving satellite aggregates the model from its covered clients every $E$ iterations, during which all participating clients within the same area perform synchronized local updates. We define ${\mathcal{I}}_{E}=\left\{{\left.{kE}\right|k=1,2,\ldots,\left\lfloor{\frac{R}{E}}\right\rfloor}\right\}$ as the set of synchronization steps of each area, where $k$ is the index of the global round. Since the devices in the same region are located nearby, we assume that they will be handed over to a new satellite parameter server at the same time according to some criteria such as distance and signal-to-noise ratio (SNR). In addition, considering partial device participation, in the $k$-th global round, the serving (nearest) satellite of area $i$ randomly selects a device set, $\mathcal{U}_{k,i}$ (with cardinality $U_{i}$), out of $\mathcal{N}_{i}$ to participate in the training process.
@@
Here, ${F_{j}}\left(\cdot\right)$ is the local loss function, defined by
@@
Let ${\mathbf{w}}_{t,j}$ be the model parameters of device $j$ at step $t\in{\mathcal{R}}$. Consider a global round of the standard FedAvg algorithm. First, when $t+1\notin\mathcal{I}_{E}$, the participating clients conduct $E$ local iterations with the latest received model parameters and local data samples. Each local iteration executes ${\mathbf{w}_{t+1,j}}={\mathbf{w}_{t,j}}-{\eta_{t}}{\nabla{F_{j}}\left({\mathbf{w}_{t,j},\varsigma_{t,j}}\right)}$, where $\eta_{t}$ is the learning rate, $\varsigma_{t,j}$ is the selected batch from device $j$ at step $t$ with $|\varsigma|$ data samples, and $\nabla{F_{j}}\left({\mathbf{w}_{t,j},\varsigma_{t,j}}\right)$ is the stochastic gradient of ${F_{j}}\left({\mathbf{w}_{t,j}}\right)$. Then, when $t+1\in\mathcal{I}_{E}$, the serving satellite aggregates the updated model parameters from these clients, conducts FedAvg by executing ${\mathbf{w}_{t+1,j}}={\frac{1}{{{N_{i}}}}\sum\limits_{j\in{{\mathcal{N}}_{i}}}\left({\mathbf{w}_{t,j}}-{\eta_{t}}{\nabla{F_{j}}\left({\mathbf{w}_{t,j},\varsigma_{t,j}}\right)}\right)}$, and broadcasts the average model back to the clients.
II-B Communication Model
The training process is conducted in consecutive global rounds until the stopping condition is reached. Without loss of generality, we focus on the global round $k$ and one ground area $i\in\mathcal{M}$ for analysis and provide detailed latency expressions as follows.
First, the selected clients conduct $E$ local iterations. For the client $j\in\mathcal{U}_{k,i}$, the computing latency of each local iteration is calculated as
$T_{j}^{\rm{local}}=\frac{|\varsigma|\Phi}{L_{j}},\;\forall j\in\mathcal{N},$ (3)
Then, clients in each area upload their models to their associated satellites. We consider orthogonal frequency-division multiple access (OFDMA) for uplink transmissions [35, 36], and each selected device per global round is assigned a dedicated sub-channel with bandwidth $W^{\rm{U}}$ [^1] In addition to digital transmission via OFDMA, analog computation methods like over-the-air aggregation become increasingly advantageous in large-scale systems, as they leverage waveform superposition to aggregate concurrent transmissions without requiring per-device orthogonalization [37, 38, 39].. Then, the uplink rate between device $j$ and its associated satellite at global round $k$ can be expressed as
${C_{k,j}^{\rm{U}}}=W^{\rm{U}}{\log_{2}}\left({1+\frac{{{P_{\rm{UE}}}{G_{\rm{UE}}}{G_{\mathrm{SAT}}}L_{k,j}^{\rm{U,PL}}L^{\mathrm{AL}}}}{{{n_{0}}W^{\rm{U}}}}}\right),$ (4)
where $P_{\rm{UE}}$ is the transmit power of ground devices, $G_{\rm{UE}}$ and $G_{\mathrm{SAT}}$ are transmitting and receiving antenna gain of device and satellite, respectively. The free-space path loss $L_{k,j}^{\rm{U,PL}}$ is defined as $L_{k,j}^{\rm{U,PL}}={\left({\frac{\lambda_{j}^{\rm{U}}}{{4\pi{d_{k,j}^{\rm{U}}}}}}\right)^{2}}$, where $\lambda_{j}^{\rm{U}}$ is the wavelength of uplink signal and $d_{k,j}^{\rm{U}}$ is the uplink transmission distance between device $j$ and its connected satellite. The additional loss $L^{\mathrm{AL}}$ is used to characterize the attenuation due to environmental factors, and $n_{0}$ is the noise power spectral density. After the serving satellite conducts FedAvg with latency $T^{\rm{agg}}$, the satellite broadcasts the averaged results to the clients, with the downlink rate between satellite and client $j$ being
${C_{k,j}^{\rm{D}}}=W^{\rm{D}}{\log_{2}}\left({1+\frac{{{P_{\rm{SAT}}}{G_{\rm{UE}}}{G_{\mathrm{SAT}}}L_{k,j}^{\mathrm{D,PL}}L^{\mathrm{AL}}}}{{{n_{0}}W^{\rm{D}}}}}\right),$ (5)
where $P_{\rm{SAT}}$ is the transmit power of satellites and $W^{\rm{D}}$ is the bandwidth that the satellite broadcasts the global model to the users. The free-space path loss of downlink transmission can be calculated by $L_{k,j}^{\rm{D,PL}}={\left({\frac{\lambda_{j}^{\rm{D}}}{{4\pi{d_{k,j}^{\rm{D}}}}}}\right)^{2}}$, where $\lambda_{j}^{\rm{D}}$ is the wavelength of downlink signal and $d_{k,j}^{\rm{D}}$ is the downlink transmission distance. In this paper, we mainly discuss users located in remote areas, where satellite links experience fewer obstacles and weaker multi-path effects compared with urban environments. This assumption is used solely for channel modeling simplification, while the proposed FedMeld framework itself remains applicable to cross-region model aggregation in other urban areas. Therefore, we ignore the impact of small-scale fading caused by the multi-path effect, being consistent with [40].
Given the data rate, the communication latency over uplink and downlink between client $j$ and its associated satellite at the global round $k$ can be expressed as $T_{k,j}^{\mathrm{U}}=\frac{q}{{C_{k,j}^{\text{U}}}}$ and $T_{k,j}^{\mathrm{D}}=\frac{q}{{C_{k,j}^{\text{D}}}}$, where $q$ denotes the model size in bits. In addition, the long transmission distance between satellites and ground clients leads that propagation latency cannot be neglected as in terrestrial networks. Thus, the propagation latency between device $j$ and its associated satellite at the global round $k$ can be expressed as $T_{k,j}^{\mathrm{P}}=\frac{d_{k,j}^{\mathrm{U}}+d_{k,j}^{\mathrm{D}}}{c}$, where $c$ is the light speed. Then, the total latency of global round $k$ can be given by
$T_{k,i}^{\mathrm{total}}=\mathop{\max}\limits_{j\in{\mathcal{U}_{k,i}}}\left\{{ET_{j}^{\mathrm{local}}+T_{k,j}^{\mathrm{U}}}+T_{k,j}^{\mathrm{D}}+T_{k,j}^{\mathrm{P}}\right\}+T_{k,i}^{\rm{agg}},$ (6)
where $T_{k,i}^{\rm{agg}}$ denotes the computation time required by the serving satellite to perform model aggregation (e.g., FedAvg) over the received client models.
Since we consider the nearest association strategy in the given satellite constellation, the serving duration of each serving satellite over a specific region remains stable [41]. This coverage duration determines the number of global training rounds the satellite can complete with the clients in a specific region, denoted by $K$. That is, $K$ is related to the characteristics of the satellite constellation, such as orbital period and satellite density.
II-C Performance Metric
For area $i$, we define an area virtual sequence ${\overline{\mathbf{w}}}_{t,i}$}} as ${{{\overline{\mathbf{w}}}_{t,i}}}={\frac{1}{{{N_{i}}}}\sum\limits_{j\in{{\mathcal{N}}_{i}}}{{{\mathbf{w}}_{t,j}}}}$. We also define a global virtual sequence as ${{\overline{\mathbf{w}}}_{t}}=\frac{1}{M}\sum\limits_{i\in{\mathcal{M}}}{{{\overline{\mathbf{w}}}_{t,i}}}$. When $t\notin\mathcal{I}_{E}$, both area and global’s sequences are inaccessible. When $t\in\mathcal{I}_{E}$, only ${\overline{\mathbf{w}}}_{t,i}$}} can be fetched for any area $i\in\mathcal{M}$. After the last training step $R$, the models can be averaged over all areas to obtain the final global model ${\overline{\mathbf{w}}}_{R}$}.